# SAGE: Sequence-level Adaptive Gradient Evolution for Generative Recommendation

### v1

@@ Abstract.
While works such as OneRec have validated the scaling laws of Large Language Models (LLMs) in recommender systems, they rely on a cumbersome separate vocabulary. This dependency prevents the model architecture from reusing native LLM vocabularies, resulting in high maintenance costs and poor scalability. In response, we aim to efficiently reuse open-source LLM architectures without constructing a separate tokenization vocabulary. Furthermore, we identify that the optimization strategy of OneRec—Gradient-Bounded Policy Optimization (GBPO)—suffers from a ”Symmetric Conservatism” problem: its static gradient boundaries structurally suppress the update momentum required for cold-start items and fail to prevent diversity collapse in high-noise environments.To address this issue, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: By combining a geometric mean importance ratio with decoupled multi-objective advantages, we eliminate token-level variance and resolve the ”Reward Collapse” problem. (2) Asymmetric Adaptive Dynamics: We construct a dynamic gradient manifold that applies a ”Boost Factor” to high-potential cold-start items to achieve super-linear updates and employs an ”Entropy-Aware Penalty” to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Recent advancements in large-scale pre-trained language models (LLMs) have introduced a transformative opportunity for the recommender system domain. These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data  (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025). Capitalizing on this potential to reshape data construction and model architecture in recommender systems holds the promise of overcoming the inherent limitations of traditional approaches, leading to enhancements in both the accuracy and diversity of personalized recommendations.
In generative recommendation systems such as OneRec, adapting an LLM to the recommendation task typically requires constructing a dedicated vocabulary to convert videos into Semantic IDs. Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges.
@@
Moreover, LLM adaptation training is mandatory. A dedicated training stage is required for the LLM to understand item tokens and absorb collaborative signals. General-purpose LLMs cannot interpret these newly defined “Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a). This includes initializing embeddings for the new tokens and freezing other parameters during training so that the LLM can learn the semantics of video tokens.
To circumvent these issues, we explore text-based models that do not require constructing a separate tokenization vocabulary, enabling efficient reuse of open-source LLM architectures. The approach is outlined as follows:
Data Construction Optimized for Large Language Models.
@@
Two-Stage Interactive Training Strategy.
SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of DeepSeek R1 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions. This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation.Additionally, high-quality samples generated in the RLHF stage will be added to the training samples. RLHF Stage: Building upon the SFT stage, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: We propose a geometric mean importance ratio combined with a decoupled multi-objective advantage estimator. This eliminates token-level variance and resolves ”Reward Collapse” in multi-objective scenarios (e.g., balancing clicks and duration). (2) Asymmetric Adaptive Dynamics: SAGE constructs a dynamic gradient manifold that is asymmetric by design. It introduces a ”Boost Factor” to allow super-linear updates for high-potential cold-start items, and an ”Entropy-Aware Penalty” that dynamically intensifies gradients for low-diversity negative samples to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
@@ 2.1. Generative Recommendation and Its Challenges
However, this paradigm introduces a dependency on a heavy vocabulary: the model must memorize the mapping between semantic tokens and specific items. The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b).
@@ 2.2. Evolution of Policy Optimization
To align large models with complex business objectives, reinforcement learning–based methods have become mainstream. GRPO  (Shao et al., 2024) eliminates the need for a Critic network by normalizing rewards within sampled groups, and it has been widely applied in reasoning tasks. However, GRPO suffers from the problem of Reward Collapse in multi-objective optimization, where samples with different performance may end up with identical advantage values after normalization, reducing the resolution of training signals. In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b).
To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization). GBPO introduces a dynamic gradient boundary derived from binary cross-entropy (BCE), effectively preventing gradient explosion caused by negative samples and improving stability in large-scale training.
@@
Therefore, we propose SAGE, a unified optimizer that is sequence-level, adaptive, and multi-objective.Token-level optimization introduces noise in Semantic ID generation. Following the logic of GSPO  (Zheng et al., 2025), SAGE elevates the importance ratio to the sequence level (geometric mean), ensuring the overall consistency of tokenized item IDs. BAPO  (Xi et al., 2025) shows that standard clipping suppresses exploration. SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-V2. To offset GBPO’s suppression of new items, SAGE incorporates the Clip-Higher philosophy from DAPO  (Yu et al., 2025), enabling super-linear updates for positive cold-start signals. To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and watch time).
@@ 3. Method
The proposed recommender system framework comprises three primary components: Data Construction, the SFT Stage, and the RLHF Stage, as depicted in Figure  2.
Figure 2. An overview of the modeling approach used in RecLLM-R1.
3.1. Data Construction
To align with the input requirements of large language models, we constructed a training dataset specifically tailored for recommender systems by transforming features into natural language descriptions. User-side Data:We transformed user profiles (e.g., age, gender, interests) and recent interaction history (e.g., sequences of clicked, viewed, liked, bookmarked, or purchased items) into chronologically ordered textual descriptions, aggregated by topic. For instance, ”Over the past 7 days, the user viewed ’Sichuan Cuisine’ videos twice and bookmarked one article titled ’Chengdu Travel Guide’.” Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions. Label Construction: Based on user engagement metrics such as browsing duration and interaction behaviors (e.g., bookmarking, sharing, commenting, following, liking), we constructed a sequence of items that reflects the user’s eventual preferences as positive training samples. This ensures the authenticity and business relevance of the labels. Prompt Construction: We concatenated the user-side and item-side textual information. The objective is to explicitly recommend the most relevant sequence of items to the user, with candidate items drawn from the provided data pool.
@@ 3.1. Data Construction
3.2. SFT Stage: Activating the Recommender Capability of the Large Language Model
@@ 3.3. RLHF Stage: Preference Alignment with Real-World User Interactions
After establishing foundational capabilities in the Supervised Fine-Tuning (SFT) stage, we introduce the RLHF phase of model. To address the cold-start and diversity issues mentioned in OneRec-V2, we propose SAGE (Sequence-level Adaptive Gradient Evolution). The core contribution of SAGE is the Asymmetric Dynamic Bounding mechanism, which adjusts the optimization trajectory based on the nature of the feedback (Positive vs. Negative) and the entropy of the generated list.
@@ 3.3.1. Adaptive Gradient Evolution
SAGE Objective Function. To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term ($1/L$):
@@
  • $\frac{1}{L}$: Length Normalization Factor. This ensures the gradient magnitude aligns with the scale of $r_{slate}$ ($\approx 1$), preventing the optimizer from being biased towards longer sequences (though $L$ is fixed here, this maintains mathematical rigor).
  • $\Phi(S,A)$: The Adaptive Dynamic Denominator, defined below.
Case A: Positive Boost for Cold-Start (Solving Suppression). When the slate receives positive feedback ($A_{SAGE}\geq 0$), standard GBPO restricts the update ratio to $r\leq 1$. This is detrimental for cold-start items where the initial probability $\pi_{old}$ is extremely low. SAGE introduces a Boost Factor:
(2) $\Phi_{pos}=\begin{cases}\frac{1}{1+\epsilon_{boost}}&\text{if }r_{slate}(\theta)>1+\epsilon_{boost}\\ 1&\text{otherwise}\end{cases}$
  • $\epsilon_{boost}$: The Positive Boost Factor (e.g., $0.2\sim 0.5$).
  • Mechanism: If the model successfully identifies a high-potential item (resulting in a high ratio $r_{slate}$), SAGE locks the denominator such that the effective gradient coefficient is maintained at $1+\epsilon_{boost}$.
  • Impact: This allows the probability of successful cold-start items to grow super-linearly, providing the necessary momentum to break through the initial exposure threshold.
Case B: Diversity-Aware Penalty (Solving Collapse). When the slate receives negative feedback ($A_{SAGE}<0$), we distinguish between ”failed exploration” and ”homogenized repetition.” We utilize the List Entropy $H(S)$ to measure the semantic diversity of the slate.
(3) $\Phi_{neg}=\max\left(1,\frac{1-r_{slate}(\theta)}{\lambda(H(S))}\right)$
@@
(4) $\begin{split}\lambda(H(S))=1+\beta\cdot\\ \tanh\left(\max(0,\bar{H}-H(S))\right)\end{split}$
  • $H(S)$: The content diversity score of the current slate (e.g., category entropy).
  • $\bar{H}$: The moving average of historical diversity.
  • $\beta$: The Diversity Temperature.
  • Mechanism:
    • Cocoon Penalty ($H(S)<\bar{H}$) If the slate is homogenized (low entropy) and rejected, $\lambda>1$. This shrinks the denominator $\Phi_{neg}$, resulting in a larger gradient (stronger penalty). This forces the model to escape the local optimum of recommending safe but boring items.
    • Exploration Tolerance ($H(S)\geq\bar{H}$) If the slate is diverse (high entropy) but rejected, $\lambda\approx 1$. The boundary reverts to the standard GBPO form, treating the error as a standard training sample without excessive punishment.

### v2

@@ Abstract.
Reinforcement learning–based preference optimization is increasingly used to align list-wise generative recommenders with complex, multi-objective user feedback, yet existing optimizers such as Gradient-Bounded Policy Optimization (GBPO) exhibit structural limitations in recommendation settings. We identify a Symmetric Conservatism failure mode in which symmetric update bounds suppress learning from rare positive signals (e.g., cold-start items), static negative-sample constraints fail to prevent diversity collapse under rejection-dominated feedback, and group-normalized multi-objective rewards lead to low-resolution training signals. To address these issues, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimizer designed for list-wise generative recommendation. SAGE introduces sequence-level signal alignment via a geometric-mean importance ratio and a decoupled multi-objective advantage estimator to reduce token-level variance and mitigate reward collapse, together with asymmetric adaptive bounding that applies positive Boost updates to successful slates and an entropy-aware penalty to discourage low-diversity failures. Experiments on Amazon Product Reviews and the large-scale RecIF-Bench demonstrate consistent improvements in top-K accuracy, cold-start recall, and diversity across both Semantic-ID and native-text action spaces, while preserving numerical stability during training. These results suggest that asymmetric, sequence-aware policy optimization provides a principled and effective framework for addressing optimization failures in generative recommendation.
@@ 1. Introduction
Recent advancements in large-scale pre-trained language models (LLMs) have introduced a transformative opportunity for the recommender system domain. These models demonstrate remarkable capabilities in semantic understanding and knowledge reasoning, and are capable of learning new tasks with limited data (Brown et al., 2020; Devlin et al., 2019; Luo et al., 2024; Li et al., 2023; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024; Li et al., 2025; Deng et al., 2025; Wei et al., 2022; Kojima et al., 2022). Capitalizing on this potential to reshape data construction and model architecture in recommender systems holds the promise of overcoming the inherent limitations of traditional approaches, leading to enhancements in both the accuracy and diversity of personalized recommendations.
In generative recommendation systems such as OneRec, adapting an LLM to the recommendation task typically requires constructing a dedicated vocabulary to convert videos into Semantic IDs. Although this design alleviates the issues of unbounded vocabulary growth and cold-start problems inherent in traditional ID-based recommendation, it still faces several practical challenges  (Zhou et al., 2025b, a; Rajput et al., 2023).
@@
Moreover, LLM adaptation training is mandatory. A dedicated training stage is required for the LLM to understand item tokens and absorb collaborative signals. General-purpose LLMs cannot interpret these newly defined ”Itemic Tokens” (video IDs); thus, vocabulary expansion and alignment must be performed during the pre-training stage  (Zhou et al., 2025b, a; Artetxe et al., 2020). This includes initializing embeddings for the new tokens and freezing other parameters during training so that the LLM can learn the semantics of video tokens.
To evaluate whether the optimization issues are intrinsic to the optimizer rather than a specific item tokenization, we consider two representative action spaces in generative recommendation: (i) Semantic-ID generation as in OneRec, and (ii) native-text item rendering without a separate item vocabulary. The approach is outlined as follows:
System instantiation for evaluation: data construction.
@@
System instantiation for evaluation: two-stage training.
SFT Stage: In the Supervised Fine-Tuning (SFT) stage, we employ a distilled version of Qwen3 to fine-tune the large language model using the constructed prompt data derived from recommender system user interactions. This process activates the model’s inherent capabilities for recommendation tasks, facilitating knowledge transfer and initial model adaptation. Additionally, high-quality samples generated in the RLHF stage will be added to the training samples. RLHF Stage: Building upon the SFT stage, we propose SAGE (Sequence-level Adaptive Gradient Evolution), a unified optimization framework tailored for list-wise generative recommendation. SAGE introduces two key innovations: (1) Sequence-level Signal Decoupling: We propose a geometric mean importance ratio combined with a decoupled multi-objective advantage estimator. This eliminates token-level variance and resolves ”Reward Collapse” in multi-objective scenarios (e.g., balancing clicks and duration). (2) Asymmetric Adaptive Dynamics: SAGE constructs a dynamic gradient manifold that is asymmetric by design. It introduces a ”Boost Factor” to allow super-linear updates for high-potential cold-start items, and an ”Entropy-Aware Penalty” that dynamically intensifies gradients for low-diversity negative samples to break information cocoons. Theoretical analysis and empirical results demonstrate that SAGE effectively unblocks cold-start traffic and sustains recommendation diversity, all while retaining the numerical stability of GBPO.
@@
Conceptual illustration of the proposed approach and its motivation for combining large language models with reinforcement learning for recommendation.
Contributions.
We identify a practical failure mode of GBPO in generative recommendation, termed Symmetric Conservatism, where symmetric conservative bounds suppress rare-item positive learning signals and can trigger diversity collapse under rejection-dominated feedback. To address this, we propose SAGE, a sequence-level, multi-objective policy optimization framework for list-wise generative recommendation. SAGE combines a sequence-level importance ratio for variance reduction, an asymmetric adaptive bounding rule with a positive boost and an entropy-aware penalty, and a decoupled normalization scheme for multi-objective rewards. We evaluate SAGE on both academic and industrial benchmarks and observe consistent gains in accuracy, cold-start discovery, and diversity, while maintaining stable training.
@@ 2.1. Generative Recommendation and Its Challenges
However, this paradigm introduces a dependency on a heavy vocabulary: the model must memorize the mapping between semantic tokens and specific items. The discrete nature of these tokens makes optimization difficult, as even slight deviations in token generation may lead to valid but irrelevant item IDs, or even entirely invalid sequences  (Deng et al., 2025; Zhou et al., 2025b; Ranzato et al., 2015).
@@ 2.2. Evolution of Policy Optimization
To align large models with complex business objectives, reinforcement learning–based methods have become mainstream. GRPO  (Shao et al., 2024) eliminates the need for a Critic network by normalizing rewards within sampled groups, and it has been widely applied in reasoning tasks. However, GRPO suffers from the problem of Reward Collapse in multi-objective optimization, where samples with different performance may end up with identical advantage values after normalization, reducing the resolution of training signals. In addition, negative feedback dominates recommendation scenarios, and directly applying standard RL often leads to gradient instability  (Zhou et al., 2025b; Zhao et al., 2017; Afsar et al., 2022; Chen et al., 2019).
To address these issues, OneRec-V2  (Zhou et al., 2025b) proposed GBPO (Gradient-Bounded Policy Optimization, OneRec-GBPO). GBPO introduces a dynamic gradient boundary derived from binary cross-entropy (BCE), effectively preventing gradient explosion caused by negative samples and improving stability in large-scale training.
@@
Therefore, we propose SAGE, a unified optimizer that is sequence-level, adaptive, and multi-objective.Token-level optimization introduces noise in Semantic ID generation. Following the logic of GSPO  (Zheng et al., 2025), SAGE elevates the importance ratio to the sequence level (geometric mean), ensuring the overall consistency of tokenized item IDs. BAPO  (Xi et al., 2025) shows that standard clipping suppresses exploration. SAGE adopts BAPO’s Entropy-Clip Rule but transforms it into a dynamic penalty term for negative samples, countering the diversity collapse observed in OneRec-GBPO. To offset GBPO’s suppression of new items, SAGE incorporates the Clip-Higher philosophy from DAPO  (Yu et al., 2025), enabling super-linear updates for positive cold-start signals. To address the reward collapse issue in GRPO, SAGE integrates the decoupled normalization strategy from GDPO  (Liu et al., 2026), achieving precise alignment across multiple objectives (e.g., clicks and Comment).
Positioning of SAGE.
Conceptually, SAGE belongs to the PPO/GBPO family of bounded policy optimization methods, but differs in two aspects that are crucial for recommendation: (i) SAGE performs sequence-level importance weighting to match list-wise feedback and reduce token-level variance; (ii) SAGE introduces an asymmetric, sample-adaptive bounding rule—boosting positive successful slates while penalizing low-entropy failures more strongly—to address cold-start suppression and diversity collapse. For multi-objective feedback, SAGE adopts a decouple-then-aggregate advantage estimator to avoid reward collapse after group normalization.
@@ 3. Method
3.1. Problem Setup
We study list-wise generative recommendation as a slate generation problem  (Ie et al., 2019). Given a user context $c$ (e.g., profile and recent behaviors), a policy $\pi_{\theta}$ generates a slate $S=(i_{1},\dots,i_{L})$ autoregressively. User feedback is provided at the slate level and can be multi-objective, denoted by $\mathbf{R}(S)=\{R_{m}(S)\}_{m=1}^{M}$. Our goal is to optimize the expected utility under mixed feedback while maintaining stable training under rejection-dominated data:
(1) $\max_{\theta}\;\mathbb{E}_{S\sim\pi_{\theta}(\cdot|c)}\left[\sum_{m=1}^{M}w_{m}R_{m}(S)\right].$
We instantiate the proposed formulation using a two-stage training pipeline (SFT followed by RLHF) on textualized recommendation data. The central algorithmic contribution is the RLHF optimizer, SAGE, which we describe in detail next and illustrate in Figure  2.
Figure 2. An overview of the modeling approach used in TextRec.
3.2. Implementation: Data Construction
To align with the input requirements of large language models, we constructed a training dataset specifically tailored for recommender systems by transforming features into natural language descriptions. User-side Data: We transformed user profiles (e.g., age, gender, interests) and recent interaction history (e.g., sequences of clicked, viewed, liked, bookmarked, or purchased items) into chronologically ordered textual descriptions, aggregated by topic. For instance, ”Over the past 7 days, the user viewed ”Sichuan Cuisine” videos twice and bookmarked one article titled ”Chengdu Travel Guide”. Item-side Data: We primarily utilized item descriptions, titles, and granular tag labels derived through content understanding, alongside statistical item features, to generate natural language descriptions. Label Construction: Based on user engagement metrics such as browsing duration and interaction behaviors (e.g., bookmarking, sharing, commenting, following, liking), we constructed a sequence of items that reflects the user’s eventual preferences as positive training samples. This ensures the authenticity and business relevance of the labels. Prompt Construction: We concatenated the user-side and item-side textual information. The objective is to explicitly recommend the most relevant sequence of items to the user, with candidate items drawn from the provided data pool.
@@ 3.2. Implementation: Data Construction
3.3. Implementation: SFT Warm-up
@@ 3.4. RLHF Stage: Preference Alignment with Real-World User Interactions
After establishing foundational capabilities in the Supervised Fine-Tuning (SFT) stage, we introduce the RLHF phase of model. To address the cold-start and diversity issues mentioned in OneRec-V2 (OneRec-GRPO), we propose SAGE (Sequence-level Adaptive Gradient Evolution). The core contribution of SAGE is the Asymmetric Dynamic Bounding mechanism, which adjusts the optimization trajectory based on the nature of the feedback (Positive vs. Negative) and the entropy of the generated list.
3.4.1. Background: GBPO in OneRec-V2
We briefly review Gradient-Bounded Policy Optimization (GBPO) proposed in the OneRec-V2 Technical Report (Zhou et al., 2025b). GBPO is a bounded policy optimization method tailored for recommendation settings with rejection-dominated feedback. It controls update magnitude by restricting the effective importance ratio to a conservative interval, which improves numerical stability compared to vanilla policy gradient. However, the symmetric and conservative design of GBPO can induce a failure mode that we term Symmetric Conservatism: (i) the upper bound suppresses learning from rare positive slates (e.g., cold-start items), and (ii) static constraints on negative updates may fail to prevent diversity collapse under repeated rejections. These limitations motivate our asymmetric and entropy-aware extension, SAGE.
@@ 3.4.2. Adaptive Gradient Evolution
SAGE as asymmetric, sample-adaptive clipping.
Let $r_{\text{slate}}(\theta)$ be the sequence-level importance ratio (defined later). SAGE optimizes a PPO-style surrogate with an asymmetric and entropy-aware clipping rule:
(2) $\mathcal{L}_{\text{SAGE}}(\theta)=\mathbb{E}\left[\mathrm{clip}_{\text{SAGE}}\!\left(r_{\text{slate}}(\theta),A_{\text{SAGE}}(S),H(S)\right)\cdot A_{\text{SAGE}}(S)\right],$
where $\mathrm{clip}_{\text{SAGE}}(\cdot)$ boosts positive updates up to $1+\epsilon_{\text{boost}}$ and applies stronger penalties to low-entropy rejected slates.
SAGE Objective Function.
To address the scale inconsistency between the geometric mean ratio (token-scale) and the sum of log-probabilities (sequence-scale), we introduce a length-normalization term $1/L$:
@@
Here, the factor $1/L$ normalizes the gradient magnitude to match the scale of $r_{\mathrm{slate}}(\theta)$ (typically around $1$), avoiding an implicit preference toward longer sequences. The term $\Phi(S,A)$ is an adaptive dynamic denominator defined below.
Case A: Positive boost for cold-start (mitigating suppression).
When the slate receives positive feedback ($A_{\mathrm{SAGE}}\geq 0$), standard GBPO effectively caps the update ratio at $r\leq 1$, which can suppress cold-start items whose initial probability under $\pi_{\mathrm{old}}$ is extremely small. We introduce a boost factor
(4) $\Phi_{\mathrm{pos}}=\begin{cases}\frac{1}{1+\epsilon_{\mathrm{boost}}},&\text{if }r_{\mathrm{slate}}(\theta)>1+\epsilon_{\mathrm{boost}},\\ 1,&\text{otherwise}.\end{cases}$
With $\epsilon_{\mathrm{boost}}\in[0.2,0.5]$, once $r_{\mathrm{slate}}(\theta)$ exceeds $1+\epsilon_{\mathrm{boost}}$, the denominator is locked so that the effective gradient coefficient is maintained at $1+\epsilon_{\mathrm{boost}}$. This enables successful cold-start items to gain probability in a super-linear manner, providing sufficient momentum to pass the initial exposure threshold.
Case B: Diversity-aware penalty (mitigating collapse).
When the slate receives negative feedback ($A_{\mathrm{SAGE}}<0$), we distinguish failed exploration from homogenized repetition by using the list entropy $H(S)$ as a proxy for semantic diversity:
(5) $\Phi_{\mathrm{neg}}=\max\!\left(1,\;\frac{1-r_{\mathrm{slate}}(\theta)}{\lambda(H(S))}\right).$
@@
(6) $\lambda(H(S))=1+\beta\cdot\tanh\!\left(\max(0,\bar{H}-H(S))\right),$
where $H(S)$ denotes the slate diversity score (e.g., category entropy), $\bar{H}$ is the moving average of historical diversity, and $\beta$ is the diversity temperature. If $H(S)<\bar{H}$, then $\lambda>1$, which reduces the denominator and yields a larger gradient magnitude, thereby imposing a stronger penalty on “cocooning” behavior. Conversely, if $H(S)\geq\bar{H}$, then $\lambda\approx 1$ and the update reverts to the standard GBPO boundary, treating the rejection as a regular training signal without excessive punishment.
@@
Algorithmic summary.
The above design can be implemented as a PPO-style on-policy update with (i) a sequence-level importance ratio, (ii) decouple-then-aggregate multi-objective advantages, and (iii) an asymmetric, entropy-aware clipping rule. For completeness, we summarize the training procedure in Algorithm LABEL:alg:sage_main.
Algorithm 1 SAGE: Sequence-level Adaptive Gradient Evolution
0: Old policy $\pi_{\theta_{\text{old}}}$, trainable policy $\pi_{\theta}$, batch of contexts $\mathcal{B}$, slate length $L$, group size $K$, objectives $\{R_{m}\}_{m=1}^{M}$, boost $\epsilon_{\text{boost}}$, diversity temperature $\beta$, clip param $\epsilon_{\text{clip}}$
0: Updated parameters $\theta$
1: Initialize running diversity target $\bar{H}$
2:for each training step do
3:  // Step 1: Sampling & Evaluation
4:  Sample batch $\mathcal{B}\sim\mathcal{D}$
5:  Generate $K$ slates $\{S_{u,k}\}$ for each user $u\in\mathcal{B}$ using $\pi_{\theta_{\text{old}}}$
6:  Compute Rewards $\{R_{m}(S_{u,k})\}$ and Diversity $H(S_{u,k})$
7:  // Step 2: Signal Decoupling (Eq. 7)
8:  for each user $u\in\mathcal{B}$ do
9:   for each objective $m$ do
10:    Normalize $R_{m}$ within group $K$: $\hat{R}_{u,k,m}\leftarrow\frac{R_{u,k,m}-\mu_{u,m}}{\sigma_{u,m}+\epsilon}$
11:   end for
12:   Aggregate: $\hat{A}_{u,k}\leftarrow\sum_{m=1}^{M}w_{m}\cdot\hat{R}_{u,k,m}$
13:  end for
14:  Batch Normalize $\hat{A}$ over entire batch $\mathcal{B}$ to get final $A_{u,k}$
15:  // Step 3: Asymmetric Adaptive Optimization
16:  for each sample $(u,k)$ in batch do
17:   Compute Sequence Ratio (Geo-Mean):
18:   $r_{u,k}\leftarrow\exp\left(\frac{1}{L}\sum_{t=1}^{L}\log\frac{\pi_{\theta}(i_{t}|c,S_{<t})}{\pi_{\theta_{\text{old}}}(i_{t}|c,S_{<t})}\right)$
19:   if $A_{u,k}\geq 0$ then
20:    Case A (Positive Boost):
21:    Set lower bound: $\epsilon_{-}\leftarrow\epsilon_{\text{clip}}$
22:    Set upper bound: $\epsilon_{+}\leftarrow\epsilon_{\text{boost}}$ {Allow super-linear update}
23:   else
24:    Case B (Diversity Penalty):
25:    Calculate regulation: $\lambda\leftarrow 1+\beta\cdot\tanh(\max(0,\bar{H}-H(S_{u,k})))$
26:    Set lower bound: $\epsilon_{-}\leftarrow\lambda-1$ {Stronger penalty if low entropy}
27:    Set upper bound: $\epsilon_{+}\leftarrow\epsilon_{\text{clip}}$
28:   end if
29:   Compute PPO Surrogate:
30:   $r_{\text{clip}}\leftarrow\text{clip}(r_{u,k},1-\epsilon_{-},1+\epsilon_{+})$
31:   $\ell_{u,k}\leftarrow\min(r_{u,k}A_{u,k},r_{\text{clip}}A_{u,k})$
32:  end for
33:  $\mathcal{L}_{\text{SAGE}}\leftarrow-\frac{1}{|\mathcal{B}|\cdot K}\sum\ell_{u,k}$
34:  Update $\theta$ by descending $\nabla_{\theta}\mathcal{L}_{\text{SAGE}}$
35:  Update $\bar{H}$ via EMA
36:end for